# Boomerang Distillation Enables Zero-Shot Model Size Interpolation

### v1

@@ 1 Introduction
Knowledge distillation has become the standard approach for producing LLM families of different sizes (Muralidharan et al., 2024). Rather than pretraining each model from scratch, practitioners often distill a pretrained teacher model into smaller student models (Hinton et al., 2015). Student models may be initialized either randomly or using parameter reduction techniques such as layer dropping (Men et al., 2024; Chen et al., 2025) or neuron pruning (Ma et al., 2023). They are then trained on large text corpora with distillation objectives, often combined with additional alignment losses such as cosine similarity or L2 distance. This paradigm is significantly more compute-efficient than independent training, reducing both FLOPs and the number of training tokens required (Muralidharan et al., 2024). However, its key limitation is that each student still requires a full training run. As a result, scaling to fine-grained model sizes remains prohibitively expensive.
@@ 2.2 Knowledge Distillation
where $\tau$ is a temperature parameter. To further align representations, we introduce a cosine distance loss (Sanh et al., 2019), which encourages the hidden states of the student across all layers to remain close to those of the teacher model. We refer to this as the alignment loss. We align the hidden states of the $i$-th layer in the student model with the hidden states produced by teacher block $\textbf{b}^{(i)}$, which corresponds to the $(\ell_{i+1}-1)$-th layer in the teacher, using a cosine distance loss:
@@ 3 Experiments
We use the same classification and generation datasets throughout the paper. We use lm-evaluation-harness (Gao et al., 2023) to evaluate all of the models and report classification accuracy on ten datasets and exact match accuracy on three generation datasets. We also compute perplexity on the WikiText dataset (Merity et al., 2017) for all models and report it in Appendix K.1. For more details on datasets, see Appendix D.
@@ 3.1 The Boomerang Distillation Phenomenon
Figure 3: Boomerang distillation emerges across model families. Shown here for Qwen3-8B, Pythia-6.9B, and Llama-3.2-3B, boomerang interpolation yields intermediate models with smooth accuracy–parameter scaling, outperforming naive layer pruning and random interpolation baselines.
@@ Results.
Figure 2 shows that boomerang distillation creates models whose size and performance interpolate smoothly (for a complete breakdown, see Appendix Figures 27 and 31). This enables us to create a full suite of intermediate models without any additional training. We show that boomerang distillation occurs when the layer-pruned, distilled student model is patched with corresponding teacher layers. In contrast, we find that the boomerang distillation phenomenon does not occur for naive layer pruning and randomly initialized distillation baselines. When we naively drop layers, there is a significant drop in classification and generation performance for models of size less than 44B inference-time parameters. However, we do not see such a dramatic drop in performance in the interpolated models created with boomerang distillation. In the randomly initialized model, there is almost no gain in performance when patching teacher layers to the distilled student. These results show that layer pruning or distillation alone is not sufficient for boomerang distillation. In Figure 3, we show that boomerang distillation also occurs in Qwen3-8B-Base, Pythia-6.9B, and Llama-3.2-3B, demonstrating that boomerang distillation is a general phenomenon in distilled LLMs that can be observed across various model sizes and families (See Appendix F for full results). We note that in the Llama model, we keep the first two layers instead of the last two layers during student initialization and patch the model starting from the first layer. This is because the first two layers of the teacher model have low cosine similarity with each other, and excluding them from the training hurts the performance of the student model and the interpolated models (see Appendix I for cosine similarity analysis).
@@
Boomerang distillation produces interpolated models that show comparable performance to the intermediate models trained via standard knowledge distillation, even outperforming them at larger sizes (Figure 4; for a per-task breakdown, see Appendix Figures 28 and 32). A key difference between the models from boomerang distillation and the standard distilled models is that we only need to train a single small student model and create interpolated models by patching teacher weights without additional training. This dramatically reduces the time and resources needed to create a family of intermediate-sized models by orders of magnitude.
@@
In Figure 4, we observe that the intermediate models at larger sizes underperform boomerang distillation models. We suspect that updating the weights of Qwen3-4B-Base on a presumably lower-quality corpus, such as The Pile, leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017), which results in a drop in performance. This is a practical problem with open-weight models because we often do not have access to the original training corpus (Jiang et al., 2023; Yang et al., 2025). Despite that, boomerang distillation can retain the benefits of the original model by patching its weights back into the student model. We show that such a drop in performance also occurs for intermediate distilled models for Llama-3.2-3B (Figure 17 in Appendix H), but not for intermediate models created by distilling Pythia-2.8B (Figure 14 in Appendix G) since it is trained on The Pile.
@@ 3.3 Effect of Knowledge Distillation
Figure 5 shows that the cross entropy with knowledge distillation loss and alignment loss (Equation 1) creates interpolated models with the lowest perplexity compared to the other loss terms (for full per-task breakdown see Appendix Figures 29 and 33). We do not see a meaningful difference in classification and generation accuracies on downstream tasks for the majority of model sizes. However, at both extremes of intermediate model size (leftmost and rightmost interpolated models), there is slight instability in performance, especially for the cases with no per-layer loss. These models correspond to patching the last few and first few teacher layers, respectively, indicating that layer-wise alignment is especially important for the first and last layers. This aligns with prior work showing that the initial and last model layers are distinct, while intermediate layers are more interchangeable  (Gromov et al., 2024; Men et al., 2024). In Appendices G and H, we demonstrate that Pythia and Llama models produce a similar ranking of loss objectives by perplexity, while also showing meaningful differences in classification accuracy.
@@ 3.5 Comparison to Layer Pruning Methods
We consider two popular layer pruning methods: Layer Collapse (LaCo) (Yang et al., 2024) and ShortGPT (Men et al., 2024). LaCo identifies blocks of layers with high cosine similarity between the outputs of the first and last layer in the block, then collapses the later layers in the block into the first one by adding their difference in parameters. ShortGPT ranks each layer in the model by its block influence, or cosine distance between the input and output activations of the layer. It then prunes the layers with the lowest block influence score. For implementation details, see Appendix L.
@@ Results.
Figure 7 shows that zero-shot interpolation via boomerang distillation results in significantly better classification and generation accuracy than layer pruning methods (for full breakdown see Appendix Figures 30 and 34). In particular, as observed in the ShortGPT paper (Men et al., 2024), the generation capabilities for the pruning approaches collapse to near zero after just a few layers removed, whereas boomerang distillation maintains higher generation accuracy for much smaller models. Boomerang distillation also smoothly interpolates in classification accuracy between the distilled student model and the teacher model, whereas the classification accuracy of all three pruning methods plateaus to near random performance for models of size around 3B parameters. We note that both of these layer pruning strategies could be used to initialize the student model in a boomerang distillation pipeline and leave exploration of different layer pruning initializations for boomerang distillation to future work.
@@ 4 Related work
Model interpolation is a key technique that combines trained models by directly interpolating their weights (Singh & Jaggi, 2020; Frankle et al., 2020; Wortsman et al., 2022). These works focus on combining weights of multiple models with additional training to improve robustness and out-of-domain generalization (Wortsman et al., 2022; Jin et al., 2023; Dang et al., 2025), create multi-task models (Ilharco et al., 2023; Yadav et al., 2023; Zhu et al., 2025), and controllable generation (Gandikota et al., 2024; Kangaslahti & Alvarez-Melis, 2024). All these works interpolate between model weights of the same size. In contrast, we interpolate between the student and the teacher model to create interpolated models of different sizes.
@@
Knowledge distillation is a popular method used to train a smaller student model to mimic the behavior of the larger teacher model with fewer parameters (Hinton et al., 2015; Sanh et al., 2019). Knowledge distillation can be used to train a smaller student model even if the teacher and the student do not share the same architecture. This has enabled researchers to distill vision models (Oquab et al., 2023), large language models (Team et al., 2024), and proprietary API-based models (Taori et al., 2023; Gudibande et al., 2024) into smaller models. Recently, knowledge distillation has been used to distill a pretrained teacher LLM into multiple smaller student LLMs of varying sizes to create a family of language models, but this approach incurs significant compute cost (Muralidharan et al., 2024; Sreenivas et al., 2024). In contrast, we use knowledge distillation to train a single student LLM using a larger teacher LLM and create interpolated models of fine-grained sizes without requiring any additional training.
@@ Acknowledgments
Appendix A Limitations
We briefly discuss the limitations of boomerang distillation.
@@ Appendix A Limitations
Boomerang distillation could also benefit from a more sophisticated student patching order. In our work, we consider two approaches to patching the student model: either starting from the first layers or the last layers. However, in some cases, this naive patching order can lead to instability in performance in the interpolated models (Appendix I). Patching the student models with teacher layers guided by the similarity of the layers could help mitigate the instability of the interpolated models.
@@ Appendix D Datasets
We utilize the same evaluation datasets throughout the paper. We use lm- evaluation-harness (Gao et al., 2023) to evaluate classification accuracy, generation exact match accuracy, and perplexity. We compute classification accuracy on $10$ tasks: ARC-easy and ARC-challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020), RACE (Lai et al., 2017), MMLU (Hendrycks et al., 2021a), and RTE (Wang et al., 2019). For generation, we report exact match accuracy on $3$ tasks: GSM8K (Cobbe et al., 2021), IFEval (Zhou et al., 2023), and MATH (Hendrycks et al., 2021b). We also compute perplexity on the WikiText dataset (Merity et al., 2017) for all experiments and report it in Appendix K.1.
@@ Results.
E.2 Impact of Training Tokens
@@ Appendix F The Boomerang Distillation Phenomenon with Qwen, Pythia, and Llama Models
F.1 Qwen3-8B-Base
@@
Figure 10 shows boomerang distillation occurs in the Qwen3-8B-Base model. Similar to Qwen3-4B-Base (§3.1), we observe a clear trend in performance as the size of the interpolated models increases. We also note that the student model created with Qwen3-8B-Base is approximately 5.6B parameters in size, which is close to the pretrained Qwen-3-4B-Base model but performs significantly worse. We suspect that the corpus used to pretrain Qwen is of a higher quality and trained on significantly more tokens compared to the distilled student model, which leads to improved out-of-the-box performance. In such cases, for a given size, we recommend choosing the model interpolated from the closest pretrained model.
F.2 Pythia-2.8B and Pythia-6.9B
@@
Figures 11 and 12 show boomerang distillation with Pythia 2.8B and Pythia 6.9B models. In both cases, we see that interpolated shows improved performance in classification accuracy, but their performance on generation tasks is nearly 0%. We also observe that the performance of the pretrained models is close to 0%, which suggests that boomerang distillation may not improve the performance of the interpolated models beyond the performance of the pretrained models.
@@
Comparing boomerang distillation to standard knowledge distillation.
@@
Comparison to layer pruning methods
@@
Comparing boomerang distillation to standard knowledge distillation.
@@
Figure 22: Per-layer cosine similarity between the output activations of the distilled student model initialized with the first two teacher layers and the teacher model, Llama 3.2 3B. After distilling a student model initialized with the first two layers (Equation 3), all student layers have high cosine similarity to their corresponding teacher block (layer ${\bm{\theta}}_{T}^{(\ell_{i})}$ shown in pink and layer ${\bm{\theta}}_{T}^{(\ell_{i+1}-1)}$ shown in yellow). Thus, patching the student model recovers smooth interpolation performance (Figure 21).
@@
Figure 23: Per-layer cosine similarity between the output activations of the distilled student model initialized with every 4th layer and the teacher model, Qwen3-4B-Base. Student layers with high cosine similarity to the outputs of their teacher blocks have predictable interpolation performance when patched in Figure 8. On the other hand, student layers with low cosine similarity see a decrease in interpolation performance when they are patched with their corresponding teacher layers.
Appendix K Additional evaluation results
@@
Comparing boomerang distillation to naive layer pruning and randomly initialized distillation.
@@
Comparing boomerang distillation to standard knowledge distillation.
@@
Comparing boomerang distillation to layer pruning methods.
@@
K.2 Classification Tasks
In this section, we report per-task classification accuracy in Figures 27-30 for experiments in §3. We find that the per-task results for all experiments align with the mean performance reported in §3.
@@
K.3 Generation Tasks
Here, we show per-task generation exact match accuracy in Figures 31-34 for experiments in §3. We find similar trends in per-task generation performance as reported for the mean generation accuracy in §3.
@@
Appendix L Pruning method details
In this section, we describe how we prune layers in the comparisons to Layer Collapse (LaCo) (Yang et al., 2024) and ShortGPT (Men et al., 2024) in Figures 7, 16, 19, 30, and 34.
@@
Table 3: Hyperparameters used to create LaCo models in Figures 7, 16, 19, 30, and 34
@@
Appendix M Use of Large Language Models

### v2

@@ 1 Introduction
Knowledge distillation has become a common approach for producing LLM families of different sizes (Muralidharan et al., 2024). Rather than pretraining each model from scratch, practitioners often distill a pretrained teacher model into smaller student models (Hinton et al., 2015). Student models may be initialized either randomly or using parameter reduction techniques such as layer dropping (Men et al., 2024; Chen et al., 2025) or neuron pruning (Ma et al., 2023). They are then trained on large text corpora with distillation objectives, often combined with additional alignment losses such as cosine similarity or L2 distance. This paradigm is significantly more compute-efficient than independent training, reducing both FLOPs and the number of training tokens required (Muralidharan et al., 2024). However, its key limitation is that each student still requires a full training run. As a result, scaling to fine-grained model sizes remains prohibitively expensive.
@@ 2.2 Knowledge Distillation
where $\tau$ is a temperature parameter. To further align representations, we introduce a cosine distance loss (Sanh et al., 2019), which encourages the hidden states of the student across all layers to remain close to those of the teacher model. We refer to this as the alignment loss. The key idea is to ensure the student layer approximates the teacher block’s output, so we can swap the teacher block back in to create interpolated models (§2.3). We align the hidden states of the $i$-th layer in the student model with the hidden states produced by teacher block $\textbf{b}^{(i)}$, which corresponds to the $(\ell_{i+1}-1)$-th layer in the teacher, using a cosine distance loss:
@@ 3 Experiments
We use the same classification and generation datasets throughout the paper. We use lm-evaluation-harness (Gao et al., 2023) to evaluate all of the models and report classification accuracy on ten datasets and exact match accuracy on three generation datasets. We also compute perplexity on the WikiText dataset (Merity et al., 2017) for all models and report it in Appendix M.1. For more details on datasets, see Appendix D.
@@ 3.1 The Boomerang Distillation Phenomenon
Figure 3: Boomerang distillation emerges across model families. Shown here for Qwen3-8B, Pythia-6.9B, and Llama-3.2-3B, boomerang distillation yields intermediate models with smooth accuracy–parameter scaling, outperforming naive layer pruning and random interpolation baselines.
@@ Results.
Figure 2 shows that boomerang distillation creates models whose size and performance interpolate smoothly (for a complete breakdown, see Appendix Figures 31 and 35). This enables us to create a full suite of intermediate models without any additional training. We show that boomerang distillation occurs when the layer-pruned, distilled student model is patched with corresponding teacher layers. In contrast, we find that the boomerang distillation phenomenon does not occur for naive layer pruning and randomly initialized distillation baselines. When we naively drop layers, there is a significant drop in classification and generation performance for models of size less than 44B inference-time parameters. However, we do not see such a dramatic drop in performance in the interpolated models created with boomerang distillation. In the randomly initialized model, there is almost no gain in performance when patching teacher layers to the distilled student. These results show that layer pruning or distillation alone is not sufficient for boomerang distillation. We also observe that boomerang distillation shows smoother interpolation in classification accuracy than in generation accuracy. This discrepancy has been observed in prior work on layer pruning: ShortGPT (Men et al., 2024) hypothesizes that errors accumulate much more in smaller pruned models than in larger ones. However, in Section 3.5, we show that boomerang distillation maintains much higher generation performance for smaller models compared to pruning methods.
In Figure 3, we show that boomerang distillation also occurs in Qwen3-8B-Base, Pythia-6.9B, and Llama-3.2-3B, demonstrating that boomerang distillation is a general phenomenon in distilled LLMs that can be observed across various model sizes and families (See Appendix F for full results). We note that in the Llama model, we keep the first two layers instead of the last two layers during student initialization and patch the model starting from the first layer. This is because the first two layers of the teacher model have low cosine similarity with each other, and excluding them from the training hurts the performance of the student model and the interpolated models (see Appendix I for cosine similarity analysis).
@@
Boomerang distillation produces interpolated models that show comparable performance to the intermediate models trained via standard knowledge distillation, even outperforming them at larger sizes (Figure 4; for a per-task breakdown, see Appendix Figures 32 and 36). A key difference between the models from boomerang distillation and the standard distilled models is that we only need to train a single small student model and create interpolated models by patching teacher weights without additional training. This dramatically reduces the time and resources needed to create a family of intermediate-sized models by orders of magnitude.
@@
In Figure 4, we observe that the intermediate models at larger sizes underperform boomerang distillation models. We suspect that updating the weights of Qwen3-4B-Base on a presumably lower-quality corpus, such as The Pile, leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017), which results in a drop in performance. This is a practical problem with open-weight models because we often do not have access to the original training corpus (Jiang et al., 2023; Yang et al., 2025). We also show that such a drop in performance also occurs for intermediate distilled models for Llama-3.2-3B (Figure 20 in Appendix H). On the other hand, we find that intermediate models created by distilling Pythia-2.8B (Figure 17 in Appendix G) perform better than interpolated models since the Pythia models are trained on The Pile. These results suggest that boomerang distillation can retain the benefits of the original model, trained on a higher-quality corpus, by patching its weights back into the student model.
@@ 3.3 Effect of Knowledge Distillation
Figure 5 shows that the cross entropy with knowledge distillation loss and alignment loss (Equation 1) creates interpolated models with the lowest perplexity compared to the other loss terms (for full per-task breakdown see Appendix Figures 33 and 37). We do not see a meaningful difference in classification and generation accuracies on downstream tasks for the majority of model sizes. However, at both extremes of intermediate model size (leftmost and rightmost interpolated models), there is slight instability in performance, especially for the cases with no per-layer loss. These models correspond to patching the last few and first few teacher layers, respectively, indicating that layer-wise alignment is especially important for the first and last layers. This aligns with prior work showing that the initial and last model layers are distinct, while intermediate layers are more interchangeable  (Gromov et al., 2024; Men et al., 2024). In Appendices G and H, we demonstrate that Pythia and Llama models produce a similar ranking of loss objectives by perplexity, while also showing meaningful differences in classification accuracy.
@@ 3.5 Comparison to Layer Pruning Methods
We consider two popular layer pruning methods: Layer Collapse (LaCo) (Yang et al., 2024) and ShortGPT (Men et al., 2024). LaCo identifies blocks of layers with high cosine similarity between the outputs of the first and last layer in the block, then collapses the later layers in the block into the first one by adding their difference in parameters. ShortGPT ranks each layer in the model by its block influence, or cosine distance between the input and output activations of the layer. It then prunes the layers with the lowest block influence score. For implementation details, see Appendix N.
@@ Results.
Figure 7 shows that zero-shot interpolation via boomerang distillation results in significantly better classification and generation accuracy than layer pruning methods (for full breakdown see Appendix Figures 34 and 38). In particular, as observed in the ShortGPT paper (Men et al., 2024), the generation capabilities for the pruning approaches collapse to near zero after just a few layers removed, whereas boomerang distillation maintains higher generation accuracy for much smaller models. Boomerang distillation also smoothly interpolates in classification accuracy between the distilled student model and the teacher model, whereas the classification accuracy of all three pruning methods plateaus to near random performance for models of size around 3B parameters. We note that both of these layer pruning strategies could be used to initialize the student model in a boomerang distillation pipeline and leave exploration of different layer pruning initializations for boomerang distillation to future work.
@@ 4 Related Work
Model interpolation is a key technique that combines trained models by directly interpolating their weights (Singh & Jaggi, 2020; Frankle et al., 2020; Wortsman et al., 2022). These works focus on combining weights of multiple models with additional training to improve robustness and out-of-domain generalization (Wortsman et al., 2022; Jin et al., 2023; Dang et al., 2025), create multi-task models (Ilharco et al., 2023; Yadav et al., 2023; Zhu et al., 2025), and controllable generation (Gandikota et al., 2024; Kangaslahti & Alvarez-Melis, 2024). All these works interpolate between model weights of the same size. In contrast, we interpolate between the student and the teacher model to create interpolated models of different sizes. Cai et al. (2025) trains the teacher model in an elastic transformer architecture along with a Gumbel Softmax-based router and interpolates model sizes using the trained router. On the other hand, boomerang distillation trains a smaller student model using a standard knowledge distillation pipeline and interpolates model sizes by patching the distilled student with teacher layers, and does not require training a specialized router.
@@
Knowledge distillation is a popular method used to train a smaller student model to mimic the behavior of the larger teacher model with fewer parameters (Hinton et al., 2015; Sanh et al., 2019). Knowledge distillation can be used to train a smaller student model even if the teacher and the student do not share the same architecture. This has enabled researchers to distill vision models (Oquab et al., 2023), large language models (Team et al., 2024), and proprietary API-based models (Taori et al., 2023; Gudibande et al., 2024) into smaller models. Recently, knowledge distillation has been used to distill a pretrained teacher LLM into multiple smaller student LLMs of varying sizes to create a family of language models, but this approach incurs significant compute cost (Muralidharan et al., 2024; Sreenivas et al., 2024). Alternatively, existing work in the vision domain  (Bai et al., 2020; Shen et al., 2020) has proposed to reduce the cost of training student models by aligning student and teacher layers in a few-shot setting, but they are limited to small-scale models and datasets and require retraining for each model size. In contrast to these approaches, we use knowledge distillation to train a single student LLM using a larger teacher LLM and create interpolated models of fine-grained sizes without requiring any additional training.
@@ Acknowledgments
Reproducibility statement
We implement all our experiments using PyTorch (Paszke et al., 2019) and HuggingFace transformers (Wolf et al., 2019) packages. We also experiment with public models available on HuggingFace Hub. We provide our code and models at https://github.com/dcml-lab/boomerang-distillation/.
@@
Appendix A Limitations and Future Work
We briefly discuss the limitations of boomerang distillation and directions for future work.
@@ Appendix A Limitations and Future Work
Boomerang distillation could also benefit from a more sophisticated initialization and student patching order. While we initialize the student model with every other layer in Section 3, we show in Appendix E.2 that initializing the student in a manner that maintains alignment allows for boomerang distillation with even smaller student models. In our work, we consider two approaches to patching the student model: either starting from the first layers or the last layers. However, in some cases, this naive patching order can lead to instability in performance in the interpolated models (Appendix  I). In Appendix  I, we analyze per-layer activation cosine similarity between all pairs of the distilled student and the base teacher layers in Llama-3.2-3B. We found that patching the student model with layers that have low cosine similarity to their teacher layers provides smoother interpolation performance. Therefore, initializing and patching the student models in a manner that is guided by the similarity of the layers could help mitigate the instability of the interpolated models.
Boomerang distillation requires the distilled student to be created via layer pruning. Combining this with other pruning strategies, such as width pruning and attention head pruning, may not work out of the box: If the teacher and student have different hidden dimensions due to width pruning, student patching cannot be applied out of the box because it would result in a mismatch in the output dimension of the hidden states. A similar obstacle occurs when pruning attention heads. Extending boomerang distillation in these settings is a promising future direction.
Finally, although we show that boomerang distillation is a phenomenon that occurs in language models, it remains to be seen whether it also occurs in other modalities, such as vision (Siméoni et al., 2025) and audio (Radford et al., 2023), that use transformer-based architectures. We leave extensions to other modalities as future work.
@@ Appendix D Datasets
We utilize the same evaluation datasets throughout the paper. We use lm- evaluation-harness (Gao et al., 2023) to evaluate classification accuracy, generation exact match accuracy, and perplexity. We compute classification accuracy on $10$ tasks: ARC-easy and ARC-challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020), RACE (Lai et al., 2017), MMLU (Hendrycks et al., 2021a), and RTE (Wang et al., 2019). For generation, we report exact match accuracy on $3$ tasks: GSM8K (Cobbe et al., 2021), IFEval (Zhou et al., 2023), and MATH (Hendrycks et al., 2021b). We also compute perplexity on the WikiText dataset (Merity et al., 2017) for all experiments and report it in Appendix M.1.
@@ Results.
E.2 Further Compressing the Distilled Model
Setup.
To test our hypothesis that the small models do not have clean interpolation behavior because of low cosine similarity, we initialize a set of small student models by manually selecting teacher layers to copy in a way that preserves alignment. To do so, we first compute the pairwise activation cosine similarity between each the output activations of each pair of teacher layers. We calculate the cosine similarity using a calibration dataset of 128 samples from The Pile (Gao et al., 2021). We then initialize the student by choosing teacher layers such that the first and last layer in each teacher block maintain high cosine similarity.
Figure 9: Boomerang distillation occurs for very small student models with cosine similarity-informed initialization Student models of size as small as 505505M parameters have relatively smooth interpolation behavior when choosing the student initialization manually to preserve alignment. This indicates that the Qwen3-4B-Base teacher model can be compressed up to 8.78.7x if the initialization is performed in a way that preserves cosine similarity between the first and last layers in each teacher block.
Results.
Figure 9 shows that by initializing the student in a way such that the teacher blocks have high similarity between the first and last layer, we can compress the student model up to 8.78.7x while preserving interpolation behavior. This validates our hypothesis from Appendix E.1 that low cosine similarity is the reason for poor interpolation performance for the naively initialized models in Figure 8. Although the student models can be compressed significantly beyond every 2nd layer, we note that these smaller models inherently produce a less granular set of interpolated models, since each student layer must be patched with its entire corresponding teacher block at once, allowing only for a maximum of M1M-1 intermediate models (where $M$ is the number of student layers).
E.3 Impact of Training Tokens
@@ Appendix F The Boomerang Distillation Phenomenon with Qwen, Pythia, and Llama Models
F.1 Qwen3-8B-Base and Qwen3-14B-Base
@@
Figure 12: Boomerang distillation with Qwen 14B creates models with smoothly interpolated size and performance.
Figures 11 and 12 show that boomerang distillation occurs in the Qwen3-8B-Base and Qwen3-14B-Base models. Similar to Qwen3-4B-Base (§3.1), we observe a clear trend in performance as the size of the interpolated models increases. We also note that the student model created with Qwen3-8B-Base is approximately 5.6B parameters in size, which is close to the pretrained Qwen-3-4B-Base model but performs significantly worse. Similarly, the student created with Qwen3-14B-Base has around 8.5B parameters and performs worse than Qwen3-8B-Base. We suspect that the corpus used to pretrain Qwen is of a higher quality and trained on significantly more tokens compared to the distilled student model, which leads to improved out-of-the-box performance. In such cases, for a given size, we recommend choosing the model interpolated from the closest pretrained model.
F.2 Pythia-2.8B, Pythia-6.9B, and Pythia 12B
@@
Figure 15: Boomerang distillation with Pythia 12B creates models with smoothly interpolated size and performance.
Figures 13, 14, and 15 show boomerang distillation with Pythia 2.8B, Pythia 6.9B, and Pythia 12B models. In all cases, we see that interpolated shows improved performance in classification accuracy, but their performance on generation tasks is nearly 0%. We also observe that the performance of the pretrained models is close to 0%, which suggests that boomerang distillation may not improve the performance of the interpolated models beyond the performance of the pretrained models.
@@
Comparison to Standard Knowledge Distillation.
@@
Comparison to Layer Pruning Methods.
@@
Comparison to Standard Knowledge Distillation.
@@
Figure 25: Per-layer cosine similarity between the output activations of the distilled student model initialized with the first two teacher layers and the teacher model, Llama 3.2 3B. After distilling a student model initialized with the first two layers (Equation 3), all student layers have high cosine similarity to their corresponding teacher block (layer ${\bm{\theta}}_{T}^{(\ell_{i})}$ shown in pink and layer ${\bm{\theta}}_{T}^{(\ell_{i+1}-1)}$ shown in yellow). Thus, student patching recovers smooth interpolation (Figure 24).
@@
Figure 26: Per-layer cosine similarity between the output activations of the distilled student model initialized with every 4th layer and the teacher model, Qwen3-4B-Base. Student layers with high cosine similarity to the outputs of their teacher blocks have predictable interpolation performance when patched in Figure 8. On the other hand, student layers with low cosine similarity see a decrease in interpolation performance when they are patched.
Appendix K Why Does Boomerang Distillation Work?
In this section, we provide an intution and proof-of-concept experiment to demonstrate why boomerang distillation works.
High-Level Intuition.
The main idea behind boomerang distillation is that we ensure through aligned initialization and distillation that not only does the student approximate the teacher model, but each student layer approximates the functionality of its corresponding teacher block. Then, patching a student layer with its corresponding teacher block produces a better approximation of the teacher model.
(a) Qwen3-4B-Base
(b) Pythia-2.8B
(c) Llama-3.2-3B
Figure 27: Last layer cosine similarity between interpolated and teacher models increases as more layers are patched acrosss models. As more student layers are patched, the interpolated models better approximate the teacher model.
Proof-of-Concept Experiment.
To show that patching a higher number of student layers creates intermediate models that better approximate the teacher, we compute the last layer cosine similarity between the activations of the interpolated and the teacher model for each number of layers patched. We use a held-out calibration set of 128 texts from The Pile (Gao et al., 2021) to compute the cosine similarity. Figure 27 demonstrates that across different models, the last layer cosine similarity between the interpolated and teacher models increases as the number of layers patched increases. This confirms our intuition that patching more layers produces interpolated models that better approximate the teacher.
Appendix L Computational Complexity
In this section, we compute the floating point operations per second (FLOPS) for boomerang distillation versus standard distillation. We compute FLOPS for $10$ iterations during training using the flops-profiler package (Li, 2023) and report the training cost computed by multiplying the average FLOPS per iteration by the total number of training iterations in Table 3. Boomerang distillation is an instance of amortized training cost, as after the initial distillation training, boomerang distillation can be used to obtain intermediate models in a zero-shot manner. As a result, boomerang distillation has equivalent computational compexity to distilling a single model of student size, and requires 14.53-19.17x less FLOPS compared to training each intermediate model independently.
Teacher Model Models FLOPS Theoretical Compute Speedup
Qwen3-4B-Base Standard distillation 8.27e20 -
Boomerang distillation 4.31e19 19.17x
Pythia-2.8B Standard distillation 4.77e20 -
Boomerang distillation 2.80e19 17.01x
Llama-3.2-3B Standard distillation 5.07e20 -
Boomerang distillation 3.49e19 14.53x
Table 3: Boomerang distillation provides significant computational speedup compared to individually distilling intermediate models. For Qwen3-4B-Base, Pythia-2.8B, and Llama-3.2-3B, we report the FLOPS required to individually distill each intermediate model versus boomerang distillation for the same number of training tokens (2.1B tokens). We can reduce FLOPs by 19.17x for Qwen, 17.01x for Pythia, and 14.53x for Llama using boomerang distillation.
Appendix M Additional Evaluation Results
@@
Comparison to Naive Layer Pruning and Random Initialization.
@@
Comparison to Standard Knowledge Distillation.
@@
Comparison to Layer Pruning Methods.
@@
M.2 Classification Tasks
In this section, we report per-task classification accuracy in Figures 31-34 for experiments in §3. We find that the per-task results for all experiments align with the mean performance reported in §3.
@@
M.3 Generation Tasks
Here, we show per-task generation exact match accuracy in Figures 35-38 for experiments in §3. We find similar trends in per-task generation performance as reported for the mean generation accuracy in §3.
@@
Appendix N Pruning Method Details
In this section, we describe how we prune layers in the comparisons to Layer Collapse (LaCo) (Yang et al., 2024) and ShortGPT (Men et al., 2024) in Figures 7, 19, 22, 34, and 38.
@@
Table 4: Hyperparameters used to create LaCo models in Figures 7, 19, 22, 34, and 38
@@
Appendix O Use of Large Language Models